a ball is thrown from a 180m high cliff> the relation h=-5t^2-5t+180 relates the time in seconds to the height in m. a) What was the intial height of the ball?
b) how long does it take the ball to reach a height of 80m? c) determine the maximum height if the ball Answer by nerdybill(7384) 
 a ball is thrown from a 180m high cliff> the relation h=-5t^2-5t+180 relates the time in seconds to the height in m.
a) What was the intial height of the ball?
this is when t=0:
h=-5t^2-5t+180
h=-5(0)^2-5(0)+180
h=180 m
.
b) how long does it take the ball to reach a height of 80m?
set h to 80 and solve for t:
h=-5t^2-5t+180
80=-5t^2-5t+180
0=-5t^2-5t+100
0=-t^2-t+20
0=t^2+t-20
0=(t+5)(t-4)
t={-5,4}
throw out the negative solution leaving:
t = 4 seconds
.
c) determine the maximum height if the ball
this occurs at the vertex when
h=-5t^2-5t+180
t = -b/(2a)
t = -(-5)/(2(-5))
t = 5/(-10)
t = -1/2 seconds
.
Note: you probably typed in a wrong equation in your original problem
probably should be something like:
h=-5t^2+5t+180